In a coal town, the coal company acts as the sole employer. Suppose the demand for coal workers is W = 5,000 - 50n, where W is the wage per month, and n is the number of coal workers hired. The supply of coal workers is given by W = 500 + 25n. If the coal company takes advantage of its monopsonist position, how many coal workers will it hire? What wage will it pay?
Answer #1

The optimum level is the equality of demand and supply.

5000 – 50n = 500 + 25n

Hence by solving,

Or, 5000 – 500 = 25n + 50n

Or, 4500 = 75n

Or, 4500/75 = n

Or, n = 60

Now, by putting this value in either demand function or supply function,

W = 5000 – 50n

Or, W = 5000 – 50 × 60

Or, W = 5000 – 3000

Or, W = 2,000

Answers:

Number of workers = 60

Wage = 2,000